Jefferson Lab HALOG ARCHIVES Logbook for December 06 - All EXCEPT Automatic Entries

  • Main INDEX

    Friday

  • 191376 12/22/06 08:01 Elena Kh. Shift Summary
  • 191374 12/22/06 07:59 Elena Kh. VDC's HVs and sp.momentums -> 0
  • 191372 12/22/06 07:14 M.Olson "Empty" target moved into beamline
  • 191371 12/22/06 07:05 Elena Kh. Beam down.
  • 191366 12/22/06 05:58 M.Olson He-3 target moved into beamline (0545)
  • 191363 12/22/06 05:34 M.Olson 20-cm dummy target into beamline at 0515
  • 191362 12/22/06 05:20 Elena Kh. spot++
  • 191359 12/22/06 05:06 Elena Kh. Hall A General Tools, Hall A Current Monitor
  • 191356 12/22/06 04:11 M.Olson Mis-labeled Q2 in beginrun/endrun comment blocks (runss 1699-1702)
  • 191350 12/22/06 03:10 atrig Trigger-Download
  • 191346 12/22/06 02:43 folts tech
  • 191345 12/22/06 02:38 M.Olson Target Screens
  • 191344 12/22/06 02:33 Elena Kh. spot++
  • 191342 12/22/06 02:20 Makis P Beam energy
  • 191341 12/22/06 01:24 atrig Trigger-Download
  • 191339 12/22/06 00:26 E.Chudakov Shift summary
  • Thursday

  • 191331 12/21/06 22:41 Mina K Right Arm VDC information
  • 191328 12/21/06 22:12 E.Chudakov Tools, beam screens
  • 191323 12/21/06 21:03 Mina k spot++
  • 191320 12/21/06 20:35 Mina K Targ controls_2
  • 191319 12/21/06 20:19 Mina K Target controls
  • 191314 12/21/06 19:21 E.Chudakov Tools and BCM Screens
  • 191311 12/21/06 18:30 Mina Katrama target screens
  • 191310 12/21/06 18:30 E.Chudakov Beam energy
  • 191300 12/21/06 16:01 Holmstrom Day Shift Summary
  • 191295 12/21/06 14:36 LeRose Right Arm up 3%/Reboot Left
  • 191289 12/21/06 13:08 A. Camsonne R VDC to 3.8 kV
  • 191287 12/21/06 13:01 Anderson VDC Efficiencies
  • 191285 12/21/06 12:24 A. Camsonne HV change right VDC to 3500 V
  • 191273 12/21/06 08:50 Anderson Plots
  • 191268 12/21/06 08:11 Holmstrom spot++
  • 191266 12/21/06 08:07 Elena Kh. Shift Summary
  • 191264 12/21/06 07:41 Elena Kh. Controlled access
  • 191263 12/21/06 07:08 Elena Kh. Cryo level on R&L Q1 dropped again
  • 191261 12/21/06 06:53 Elena Kh. Hall A General Tools
  • 191260 12/21/06 06:50 M.Olson Mislabeled start run entries (runs 1665-1668)
  • 191255 12/21/06 05:26 M.Olson Target Screens
  • 191248 12/21/06 03:08 M.Olson Argon gas supply bottle changed (automatically)
  • 191235 12/21/06 02:30 Elena Kh. spot++
  • 191227 12/21/06 02:19 Elena Kh. spot++
  • 191215 12/21/06 02:06 Elena Kh. controlled access
  • 191213 12/21/06 00:06 folts tech Q3
  • 191212 12/21/06 00:03 e04018 Shift Summary
  • Wednesday

  • 191210 12/20/06 23:47 atrig Trigger-Download
  • 191206 12/20/06 20:15 E. Cisbani Target Status
  • 191204 12/20/06 18:27 F. Garibaldi Trips on Q1 and dipole
  • 191200 12/20/06 17:59 E. Cisbani Coda restarted
  • 191192 12/20/06 16:03 Holmstrom Day Shift - Summary
  • 191182 12/20/06 14:42 atrig Trigger-Download
  • 191168 12/20/06 12:21 Holmstrom Left HRS IOC
  • 191165 12/20/06 11:55 Anderson VDC Efficiencies
  • 191151 12/20/06 08:06 Elena Kh. Shift Summary
  • 191150 12/20/06 07:56 Elena Kh. Hall A General Tools, Hall A Current Monitor
  • 191147 12/20/06 07:45 M.Olson VDC efficiency plots (run 1624)
  • 191144 12/20/06 07:12 M.Olson Hall A tools and current draw of right (top) VDC
  • 191133 12/20/06 05:59 M.Olson VDC efficiency plots (run 1618)
  • 191124 12/20/06 03:32 M.Olson VDC efficiency plots (run 1614)
  • 191121 12/20/06 02:34 M.Olson Target screens
  • 191116 12/20/06 01:53 Elena Kh. VDC eff. from run #1610
  • 191113 12/20/06 01:36 M.Olson Hall A Tools for run 1610
  • 191112 12/20/06 01:35 M.Olson spot++ for run 1610
  • 191098 12/20/06 01:08 Elena Kh. new R S1 HVs
  • 191097 12/20/06 01:03 atrig Trigger-Download
  • 191096 12/20/06 00:57 Elena Kh. R S1 HV
  • 191094 12/20/06 00:43 gomez Left VDCs Gas Flow Adjustment
  • 191093 12/20/06 00:42 jtewing techs
  • Tuesday

  • 191091 12/19/06 22:31 e04018 right arm dipole power supply not responding
  • 191086 12/19/06 21:16 A. Camsonne Left arm VDC Ok Right VDC recovering
  • 191085 12/19/06 20:51 E. Cisbani Target moved to He3 20cm
  • 191080 12/19/06 20:06 A. Camsonne Harp runs summary
  • 191077 12/19/06 19:43 A. Camsonne Harp Scan (1,-1)
  • 191076 12/19/06 19:41 A. Camsonne VDC efficiency run 1597
  • 191073 12/19/06 19:24 A. Camsonne Harp Scan (1,0)
  • 191070 12/19/06 19:14 A. Camsonne Harp Scan (1,1)
  • 191067 12/19/06 19:03 A. Camsonne Harp Scan (0,1)
  • 191064 12/19/06 18:48 A. Camsonne Harp Scan (0,-1)
  • 191061 12/19/06 18:39 A. Camsonne Harp Scan (-1,-1)
  • 191058 12/19/06 18:33 A. Camsonne Harp Scan (-1,0)
  • 191055 12/19/06 18:14 A. Camsonne Harp Scan (0,0)
  • 191051 12/19/06 17:56 E. Cisbani Targen on optics (carbon foils)
  • 191049 12/19/06 17:46 A. Camsonne Left VDC recovering too run 1589
  • 191048 12/19/06 17:22 A. Camsonne Readout of flow for right arm
  • 191047 12/19/06 17:14 A. Camsonne Correction Right Arm VDC plot for run 1586
  • 191045 12/19/06 17:10 A. Camsonne Left arm VDC efficiency run 1587
  • 191043 12/19/06 16:58 A. Camsonne Right arm VDC recovering ! Run 1587
  • 191036 12/19/06 16:01 Holmstrom Day Shift Summary
  • 191031 12/19/06 15:50 atrig Trigger-Download
  • 191028 12/19/06 15:31 folts detector gas
  • 191026 12/19/06 14:31 Holmstrom 1582 - Left arm
  • 191025 12/19/06 14:29 Holmstrom 1582 - Rt Arm VDC eff
  • 191022 12/19/06 13:32 Holmstrom Left Arm VDC eff. 1581
  • 191021 12/19/06 13:31 Holmstrom Right Arm VDC 1581
  • 191017 12/19/06 12:46 Holmstrom 1579 Right arm VDCs
  • 191016 12/19/06 12:44 Holmstrom 1579 VDC eff
  • 191012 12/19/06 12:24 Holmstrom Right arm VDC eff
  • 191011 12/19/06 11:13 Holmstrom VDC Eff plots
  • 191002 12/19/06 09:23 Holmstrom Right arm eff.
  • 191001 12/19/06 09:09 Holmstrom VDC eff problem
  • 190998 12/19/06 08:07 Elena Kh. Shift Summary
  • 190995 12/19/06 06:14 Elena Kh. N2 level in L Q2
  • 190994 12/19/06 06:07 M.Olson Target Screens
  • 190993 12/19/06 05:40 Elena Kh. Test run for VDC eff.: Right arm
  • 190992 12/19/06 05:38 Elena Kh. Test run for VDC eff. : Left arm
  • 190990 12/19/06 04:56 Elena Kh. Hall A General Tools
  • 190988 12/19/06 04:35 M.Olson Attempted Harp scans
  • 190981 12/19/06 03:17 Elena Kh. More on VDC problems
  • 190980 12/19/06 02:50 gomez VDC Efficiency Problems!
  • 190979 12/19/06 02:42 M.Olson Update on VDC problem - currently down ~4 hrs
  • 190974 12/19/06 00:26 E.Chudakov Shift summary
  • Monday

  • 190970 12/18/06 23:41 E.Chudakov Problems, run stopped
  • 190968 12/18/06 22:40 E.Chudakov Beam is back after a 70 min break
  • 190965 12/18/06 21:36 E.Chudakov Harp scan not done
  • 190959 12/18/06 17:22 E.Chudakov HRSL moved to kin8: 30.65 deg, 2.803 GeV/c
  • 190952 12/18/06 15:59 jpchen shift summary
  • 190941 12/18/06 14:51 jpchen BCM calibration constants
  • 190933 12/18/06 13:51 jpchen TS11 stuck, rebooted and restarted coda
  • 190931 12/18/06 13:50 jpchen BCM reading went wrong
  • 190927 12/18/06 13:12 jpchen half-wave plate insertion
  • 190921 12/18/06 12:34 jpchen beam back, left dipole set
  • 190920 12/18/06 11:19 jpchen Mott from 10:00, Left dipole
  • 190916 12/18/06 08:01 E. Cisbani Owl Shift Summary
  • 190912 12/18/06 07:12 M.Olson spot++ for run 1543
  • 190911 12/18/06 05:55 E. Cisbani Genelar Tools and Beam Chart
  • 190908 12/18/06 04:31 M.Olson Reboot right NMR during run 1542
  • 190905 12/18/06 02:59 M.Olson Target Screens
  • 190904 12/18/06 02:44 M.Olson Correction to item (2.) in previous half-wave plate halog entry
  • 190903 12/18/06 02:14 M.Olson More on half-wave plate status
  • 190900 12/18/06 01:14 M.Olson Half-wave plate changes
  • 190899 12/18/06 00:20 E. Cisbani spot++ run 1540
  • 190898 12/18/06 00:08 e04018 shift summary
  • Sunday

  • 190894 12/17/06 23:23 Ramesh Hall A tools
  • 190891 12/17/06 20:50 Ramesh Target status
  • 190884 12/17/06 16:01 E. Cisbani Day shift summary
  • 190880 12/17/06 13:30 S. Frullani Target status
  • 190877 12/17/06 12:03 E. Cisbani Hall A Current Monitor
  • 190876 12/17/06 11:19 E. Cisbani HallA Tool screen
  • 190875 12/17/06 11:17 E. Cisbani spot++ run 1533
  • 190873 12/17/06 11:04 E. Cisbani 1/2 Wave Plane Now IN
  • 190868 12/17/06 08:00 Elena Kh. Shift Summary
  • 190863 12/17/06 04:43 M.Olson spot++ (run 1529) - Looks more like 1.5 x 2 as opposed to 2 x 2
  • 190862 12/17/06 04:31 Elena Kh. Hall A General Tools, Current Monitor
  • 190861 12/17/06 03:25 M.Olson Target screens
  • 190858 12/17/06 01:02 Elena Kh. Rebooted ioc for right magnet and R NMR
  • Saturday

  • 190854 12/16/06 23:49 Ramesh Shift Summary
  • 190851 12/16/06 21:42 Ramesh Hall A tools
  • 190848 12/16/06 20:12 E. Cisbani Target screens snapshot
  • 190847 12/16/06 19:59 Ramesh Subed Left hrs magnet readback frozen again
  • 190842 12/16/06 16:50 Ramesh Subed IOC for left hrs magnets rebooted
  • 190841 12/16/06 16:18 M. Petratos Shift Summary
  • 190833 12/16/06 10:29 Salvatore anomaly in target vacuum
  • 190830 12/16/06 08:04 M.Olson (cryotarget) Target vacuum readback becoming noisy
  • 190828 12/16/06 08:01 R. Michaels owl summary
  • 190827 12/16/06 07:28 M.Olson Second try at the Kinematic 8 run summary
  • 190826 12/16/06 06:51 M.Olson Summary of He3 production runs thusfar on kinematic point 8
  • 190825 12/16/06 06:28 M.Olson Lost communication from RHRS, rebooted ioc and NMR
  • 190820 12/16/06 04:14 M.Olson Target Screens
  • 190819 12/16/06 04:05 M.Olson (Tech) N2 level for CLL72Q110 (Right) reads zero on Hall A Tools
  • 190818 12/16/06 03:43 M.Olson spot++ for run 1517
  • 190813 12/16/06 00:22 M.Olson spot++ for run 1515
  • Friday

  • 190811 12/15/06 23:57 Ramesh Subed Shift Summary
  • 190808 12/15/06 22:16 Salvatore target status
  • 190807 12/15/06 22:08 A. Camsonne Timing offset changed for R S2
  • 190804 12/15/06 20:35 Adaq spot++spot++
  • 190803 12/15/06 19:48 Jack iocha16 is hung up
  • 190797 12/15/06 15:47 E. Cisbani run summary
  • 190796 12/15/06 15:46 E. Cisbani JT valve opened 1 step
  • 190795 12/15/06 15:02 E. Cisbani spot++ for run 1511
  • 190791 12/15/06 13:19 E. Cisbani spot++ for run 1510
  • 190789 12/15/06 11:16 E. Cisbani spot++ for run 1509
  • 190784 12/15/06 08:28 FANSLER L Q2 N2 / TECHS
  • 190782 12/15/06 08:03 M.Olson Owl shift summary
  • 190781 12/15/06 07:35 itra Shabest Target Screen
  • 190780 12/15/06 07:23 M.Olson Beamline LeCroy HV Alarm
  • 190779 12/15/06 07:00 Mitra Shabes Run Summary
  • 190778 12/15/06 06:52 M.Olson spot++ for run 1507
  • 190777 12/15/06 06:51 M.Olson Hall A tools after LHRS ioc reboot
  • 190775 12/15/06 06:50 M.Olson Lost communication with LHRS again, rebooted ioc
  • 190773 12/15/06 05:26 M.Olson spot++ for run 1506
  • 190772 12/15/06 05:25 M.Olson Hall A Tools and Stable Beam Current Profile
  • 190769 12/15/06 03:34 M.Olson Hall A Tools after LHRS ioc reboot
  • 190768 12/15/06 03:31 Mitra Shabes Target Screen
  • 190767 12/15/06 03:26 M.Olson spot++ for run 1505
  • 190765 12/15/06 03:11 M.Olson Lost communication from LHRS, rebooted ioc
  • 190763 12/15/06 02:47 M.Olson Asked for slight (2uA) decrease in current to stabilize target control
  • 190762 12/15/06 02:43 M. Olson Closed JT valve by 0.5% to bring CHL flow below 20g/s
  • 190761 12/15/06 02:10 M.Olson spot++ for run 1504
  • 190758 12/15/06 00:13 M.Olson Hall A Tools for run 1503
  • 190757 12/15/06 00:09 M.Olson spot++ for run 1503
  • Thursday

  • 190754 12/14/06 23:59 Ramesh Subed Shift Summary
  • 190752 12/14/06 22:15 Ramesh Hall A tools and current monitor
  • 190747 12/14/06 17:55 Ramesh Target Status
  • 190744 12/14/06 16:16 F. Garibaldi Shift Summary
  • 190738 12/14/06 15:12 Makis P Limited to 85uA by target/Maxed CHL flow
  • 190737 12/14/06 13:17 Makis P New Kinematics (He3 #8) Q2=68fm-2 Forward
  • 190734 12/14/06 12:47 Makis P 85-90 uA/Opened JT Valve one step
  • 190733 12/14/06 12:27 Elena Kh. spot++
  • 190729 12/14/06 12:07 R. Michaels scaler.map fix
  • 190726 12/14/06 11:55 Makis P He3 target back in beam
  • 190723 12/14/06 11:40 Elena Kh. Ped. sub. on
  • 190721 12/14/06 11:23 Elena Kh. Ped. sub. off
  • 190717 12/14/06 09:43 gunning raster
  • 190716 12/14/06 09:08 atrig Trigger-Download
  • 190713 12/14/06 08:01 M.Olson Owl Shift Summary
  • 190712 12/14/06 07:51 Mitra Shabes Owl Shift run summary
  • 190711 12/14/06 07:34 Mitra Shabes Target Screen
  • 190710 12/14/06 07:31 M.Olson Triggers for run 1489 still look OK despite incorrect prescales
  • 190709 12/14/06 07:09 M.Olson Incorrect prescales for run 1489
  • 190706 12/14/06 06:39 M.Olson Hall A Tools, run 1489
  • 190705 12/14/06 06:38 M.Olson spot++ for run 1489
  • 190703 12/14/06 06:26 Mitra Shabes Moved target to 3He
  • 190702 12/14/06 06:19 M.Olson spot++ after raster power supply swap (run 1488)
  • 190699 12/14/06 06:07 R. Michaels raster repair
  • 190696 12/14/06 04:46 Mitra Shabes Target at empty position
  • 190695 12/14/06 04:46 M.Olson Update on raster issue
  • 190694 12/14/06 04:00 M.Olson Made contact with Run Coordinator
  • 190693 12/14/06 03:40 M.Olson Paged and left voicemail for run coordinator
  • 190692 12/14/06 03:28 M.Olson Spoke with on-duty technician, paging Bob Michaels
  • 190691 12/14/06 03:20 M. Olson Lost vertical raster, trying to contact on-duty tech
  • 190689 12/14/06 02:14 Mitra Shabes ESR_Flow_(HallA) fluctuation
  • 190688 12/14/06 01:40 Mitra Shabes Target Screen
  • 190687 12/14/06 01:26 M.Olson Spot++ for run 1486
  • 190684 12/14/06 00:45 M.Olson Hall A Tools and Beam Current Profile
  • 190683 12/14/06 00:32 M.Olson spot++ for run 1485
  • 190681 12/14/06 00:07 Ramesh Subed Shift Summary
  • Wednesday

  • 190672 12/13/06 18:47 S. Frullani snapshot target
  • 190669 12/13/06 18:03 Adaq Hall A tools and Current Strip chart
  • 190668 12/13/06 17:01 Ramesh Subed Hall A BCM coefficients changed
  • 190664 12/13/06 16:04 LeRose Shift Summary
  • 190663 12/13/06 15:46 LeRose Target BPM Strip Tool
  • 190662 12/13/06 14:00 atrig Trigger-Download
  • 190661 12/13/06 13:32 Michaels/Cam Trigger 6 issues
  • 190659 12/13/06 13:21 A. Camsonne Right arm aluminium VDC covers removed
  • 190658 12/13/06 12:56 LeRose Reboot Left Arm Controls
  • 190657 12/13/06 12:50 R. Michaels T6 modifications
  • 190656 12/13/06 12:44 atrig Trigger-Download
  • 190655 12/13/06 12:43 atrig Trigger-Download
  • 190654 12/13/06 12:42 atrig Trigger-Download
  • 190653 12/13/06 12:31 atrig Trigger-Download
  • 190652 12/13/06 12:29 atrig Trigger-Download
  • 190651 12/13/06 12:15 atrig Trigger-Download
  • 190650 12/13/06 11:31 atrig Trigger-Download
  • 190649 12/13/06 11:30 R. Michaels T4 cerenkov threshold
  • 190648 12/13/06 11:09 A. Camsonne Aluminium protection for VDC put on
  • 190647 12/13/06 10:56 LeRose Hall A Tools
  • 190646 12/13/06 10:55 LeRose NMR reboot successful
  • 190645 12/13/06 10:15 LeRose Beam Off/Controlled Access
  • 190643 12/13/06 10:12 LeRose Spot++
  • 190642 12/13/06 10:03 LeRose Right NMR not updating
  • 190641 12/13/06 08:14 R. Michaels scaler.map updated in onlana
  • 190639 12/13/06 07:57 M.Olson Owl shift summary
  • 190638 12/13/06 07:34 M.Olson spot++ for run 1478
  • 190637 12/13/06 07:33 M.Olson Hall A Tools (run 1478)
  • 190636 12/13/06 07:32 Mitra Shabes Target Screen
  • 190635 12/13/06 07:25 M. Olson Problems getting NMR (L) to lock
  • 190633 12/13/06 05:44 M.Olson Lost LHRS ioc, rebooting
  • 190631 12/13/06 04:18 M.Olson Spot++ for run 1477
  • 190628 12/13/06 02:41 M. Olson spot++ for run 1476
  • 190627 12/13/06 02:40 M. Olson HAll A Tools and Current Profile (run 1476)
  • 190626 12/13/06 02:37 Mitra Shabes Target Screen
  • 190623 12/13/06 00:46 M.Olson Spot++ for run 1475
  • 190622 12/13/06 00:36 M.Olson Hall A Tools and Beam Current Profile (Run 1475)
  • 190619 12/13/06 00:04 Mina Katrama Shift Summary
  • Tuesday

  • 190617 12/12/06 22:45 Mina K General tools screen
  • 190616 12/12/06 22:43 Makis P 1469 End Run entry and dual 1470 End of Run entries
  • 190615 12/12/06 22:30 A. Camsonne BCM calibration constant for scalers updated
  • 190610 12/12/06 21:35 Makis P Target charts 9:30pm Run 1472
  • 190609 12/12/06 21:30 Mina K spot++_2
  • 190608 12/12/06 21:27 Mina K spot++
  • 190605 12/12/06 20:08 Mina K General Tools screen
  • 190604 12/12/06 19:14 M.Iodice runs 1469 and 1470 END_OF_RUN files
  • 190600 12/12/06 18:35 Mina K Current monitor
  • 190597 12/12/06 16:29 Mina K spot++
  • 190594 12/12/06 16:11 R. Michaels T4 seems wrong
  • 190591 12/12/06 16:06 Higinbotham Shift Summary
  • 190589 12/12/06 15:53 S. Frullani General Tools & Beam Current
  • 190580 12/12/06 08:02 M.Olson spot++ for run 1463
  • 190578 12/12/06 08:00 M. Olson Owl Shift Summary
  • 190575 12/12/06 07:35 Mitra Shabes Target Screen
  • 190574 12/12/06 07:10 M. Olson Beam Current Profile over last two hours
  • 190573 12/12/06 06:00 M.Olson Fixed beam half-wave status for run 1461
  • 190570 12/12/06 05:41 M.Olson Beam half-wave status error in endruns (1454-1461)
  • 190569 12/12/06 05:24 Mitra Shabes Target Screen
  • 190568 12/12/06 05:07 Mitra Shabes Reviewed online analysis for runs 1457 & 1460.
  • 190567 12/12/06 05:03 M.Olson Beam current profile for last two hours
  • 190566 12/12/06 05:00 M.Olson Hall A General Tools (run 1461)
  • 190565 12/12/06 04:51 M. Olson Spot++ for run 1461
  • 190564 12/12/06 04:36 M.Olson (DAQ expert please read) Trigger numbers for run 1461
  • 190561 12/12/06 03:54 M.Olson Lost coda during run 1459, restarted with run 1460
  • 190559 12/12/06 03:36 M.Olson Hall A Tools after RHRS IOC Reboot
  • 190558 12/12/06 03:19 M.Olson RHRS Readbacks appear frozen restarted IOC 's
  • 190555 12/12/06 02:10 M. Olson Hall A General Tools
  • 190554 12/12/06 02:09 M.Olson Spot++ for run 1458
  • 190553 12/12/06 01:59 M.Olson Beam Current Profile for last two hours
  • 190550 12/12/06 01:52 A. Camsonne Pion rejector
  • 190549 12/12/06 01:18 Mitra Shabes Target Screen
  • 190548 12/12/06 00:38 M.Olson Spot++ for run 1457
  • Monday

  • 190546 12/11/06 23:59 Mina Katrama Mina Katramatou
  • 190543 12/11/06 23:45 Mina K beam current monitor
  • 190542 12/11/06 23:03 Mina Katrama General Tools screen
  • 190541 12/11/06 23:01 Makis Petrat Target Screens
  • 190540 12/11/06 21:55 Mina k Beam Current Monitor
  • 190539 12/11/06 21:52 Mina K spot++
  • 190537 12/11/06 21:49 Mina K Beam back
  • 190536 12/11/06 20:07 Makis Petrat target screens
  • 190535 12/11/06 20:02 Makis Petrat target screens
  • 190534 12/11/06 17:37 Mina Katrama Beam Off - temp. alarm
  • 190532 12/11/06 16:36 Mina Katrama Current monitor screen
  • 190531 12/11/06 16:29 Mina Katrama spot++
  • 190530 12/11/06 16:02 Elena Kh. Shift Summary
  • 190526 12/11/06 15:22 A. Camsonne Pedestal subtraction back on
  • 190524 12/11/06 14:24 Elena Kh. run #1454
  • 190522 12/11/06 14:18 Elena Kh. Target screens
  • 190519 12/11/06 13:34 A. Camsonne Turned pedestal subtraction to check for low amplitudes
  • 190518 12/11/06 13:26 Elena Kh. He3 target
  • 190517 12/11/06 08:55 Elena Kh. beam study
  • 190514 12/11/06 08:01 M .Olson Owl Shift Summary
  • 190513 12/11/06 07:51 Mitra Shabes Target Screen
  • 190512 12/11/06 07:42 M.Olson spot++ for run 1452
  • 190510 12/11/06 07:35 M.Olson Beam back
  • 190508 12/11/06 07:33 M.Olson MCC called, doing some tuning.
  • 190507 12/11/06 07:09 M.Olson spot++ for run 1451
  • 190504 12/11/06 06:35 M.Olson Beam current profile over last two hours
  • 190503 12/11/06 05:25 M.Olson Spot++ for run 1450
  • 190500 12/11/06 04:07 Mitra Shabes Target Screen
  • 190495 12/11/06 02:18 M. Olson spot++ for run 1447
  • 190494 12/11/06 02:17 M. Olson Hall A General Tools Screen Grab
  • 190491 12/11/06 00:05 Mina Katrama Shift Summary
  • Sunday

  • 190487 12/10/06 23:04 Mina Katrama General tools screen
  • 190484 12/10/06 21:15 Makis P target screens
  • 190483 12/10/06 20:06 Mina Katrama spot++
  • 190475 12/10/06 19:27 atrig Trigger-Download
  • 190474 12/10/06 18:20 M.Olson Fan power deposition not included in total loop power
  • 190470 12/10/06 17:03 Mina Katrama Beam Current Monitor
  • 190469 12/10/06 16:49 Mina katrama spot++
  • 190467 12/10/06 16:03 Kumbartzki Day shift Summary
  • 190464 12/10/06 15:41 Mitra Shabes Target Screen
  • 190461 12/10/06 13:17 Mitra Shabes Target screen
  • 190460 12/10/06 13:14 Mitra Shabes Target moved to 3He
  • 190457 12/10/06 12:46 Mitra Shabes Target screen
  • 190456 12/10/06 12:46 Mitra Shabes Target Moved to Al dummy
  • 190451 12/10/06 08:57 kumbartzki spot of run 1434
  • 190449 12/10/06 08:15 Elena Kh. Shift Summary
  • 190446 12/10/06 08:14 M. Olson Changes to target alarm settings
  • 190445 12/10/06 08:07 Elena Kh. current monitor
  • 190444 12/10/06 07:53 Elena Kh. spot++
  • 190442 12/10/06 06:11 Elena Kh. Hall A General Tools
  • 190441 12/10/06 05:05 M. Olson MCC estimate for beam restoration = two hours
  • 190440 12/10/06 05:04 M. Olson Target restored to nominal conditions
  • 190439 12/10/06 04:05 M. Olson Rebooted Target IOC's
  • 190438 12/10/06 03:32 M. Olson Target update
  • 190437 12/10/06 03:13 Elena Kh. runs #1427-1431
  • 190434 12/10/06 01:38 M. Olson Cryo reports ready to start cooling
  • 190433 12/10/06 01:37 M. Olson Initial target recovery instructions from Mike Seely
  • 190432 12/10/06 01:21 fansler techs/power loss
  • 190431 12/10/06 00:49 M. Olson Current Target Status
  • 190429 12/10/06 00:12 Mina Katrama Shift summary
  • 190428 12/10/06 00:07 Mitra Shabes Target
  • Saturday

  • 190427 12/09/06 23:17 Mina Katrama Power glitch
  • 190421 12/09/06 20:29 M. Shabestar Target Screen
  • 190418 12/09/06 18:55 Mina katrama Spot++
  • 190417 12/09/06 18:53 mina katrama beam current monitor
  • 190416 12/09/06 17:49 Mina Katrama beam back
  • 190413 12/09/06 15:59 kumbartzki day shift summary
  • 190407 12/09/06 13:12 Kumbartzki spot
  • 190406 12/09/06 12:29 kumbartzki current chart
  • 190396 12/09/06 07:58 Adaq Left Arm HV
  • 190395 12/09/06 07:55 suleiman Owl Shift Summary
  • 190394 12/09/06 07:52 Adaq Current Chart
  • 190391 12/09/06 06:49 Elena Kh. spot++
  • 190390 12/09/06 06:47 Elena Kh. Target screens
  • 190385 12/09/06 04:33 Adaq Rates during run 1418 at 60 uA
  • 190384 12/09/06 04:05 Adaq Beam Current Chart
  • 190383 12/09/06 03:47 Elena Kh. spot++
  • 190382 12/09/06 03:37 Elena Kh. target vacuum screen
  • 190377 12/09/06 02:44 Elena Kh. Target screens
  • 190374 12/09/06 01:33 Adaq spot++
  • 190373 12/09/06 01:33 Adaq Right Arm HV
  • 190372 12/09/06 01:32 Adaq Hall A General Tools
  • 190371 12/09/06 01:31 Adaq Tried to go to 70 uA but did go well
  • 190370 12/09/06 01:10 Elena Kh. Target screens
  • 190369 12/09/06 01:08 Elena Kh. runs #1409-1414
  • 190353 12/09/06 00:24 E.Chudakov Shift summary
  • Friday

  • 190340 12/08/06 20:26 E.Chudakov Raster and beam position
  • 190338 12/08/06 20:17 E.Chudakov Beam current calibration
  • 190331 12/08/06 17:23 R.Feuerbach corrected decoding of R.s1
  • 190330 12/08/06 16:17 fansler left dipole He 80% max read back
  • 190329 12/08/06 16:06 jpchen day shift summary
  • 190327 12/08/06 15:45 jpchen MCC inform: half-wave plate removed
  • 190326 12/08/06 15:38 A. Camsonne Temporary fix S1 display
  • 190325 12/08/06 15:27 Saha FC2/OLO2 calib data
  • 190324 12/08/06 13:29 jpchen MCC is still struggling with current calibration at high current
  • 190323 12/08/06 10:55 Kumbartzki empty target
  • 190322 12/08/06 10:47 jpchen Hall A tools screen: normal
  • 190321 12/08/06 10:46 jpchen BCM plot. Beam tuning for the last near two hours
  • 190316 12/08/06 07:58 suleiman Owl Shift Summary
  • 190315 12/08/06 07:08 A. Camsonne Data for Right S1 PMT 6 not being decoded correctly
  • 190314 12/08/06 06:28 Adaq BeaM Current Screen
  • 190313 12/08/06 06:27 Elena Kh. spot++
  • 190310 12/08/06 05:47 Elena Kh. spot++
  • 190303 12/08/06 03:29 Adaq Beam Current the last 2 hours
  • 190302 12/08/06 02:54 Adaq Energy Lock changed to Hall C line
  • 190299 12/08/06 02:29 Adaq Target Screens
  • 190298 12/08/06 02:26 suleiman Right S1 L-PMT 5 has no ADC
  • 190297 12/08/06 02:15 Adaq Right Arm HV
  • 190296 12/08/06 01:45 Adaq spot++
  • 190295 12/08/06 01:44 Adaq Hall A General Tools Screen
  • 190292 12/08/06 00:24 A. Camsonne T6 trigger coincidence of T2 and T3
  • 190291 12/08/06 00:20 A. Camsonne Right arm S1 Left paddle 5 ( index 4 ) moved to ADC slot 25 channel 48
  • 190290 12/08/06 00:03 Seonho Choi Swing shift summary (Dec. 7, 2006)
  • 190288 12/08/06 00:01 Seonho Choi Beam unstable beyond 50uA
  • Thursday

  • 190287 12/07/06 23:58 Kumbartzki target screen
  • 190284 12/07/06 19:08 atrig Trigger-Download
  • 190272 12/07/06 16:46 atrig Trigger-Download
  • 190269 12/07/06 16:09 Lindgren Hall A Tools end of shift
  • 190267 12/07/06 16:06 Lindgren Shift Summary
  • 190264 12/07/06 13:27 Lindgren HRS
  • 190262 12/07/06 11:43 Lindgren Noon update
  • 190254 12/07/06 10:26 Lindgren Charts
  • 190253 12/07/06 10:23 Lindgren Charts
  • 190251 12/07/06 10:09 Lindgren Cryotarget Controls
  • 190249 12/07/06 10:03 R.A. Lindgre Beam off/Hall A Controlled Access
  • 190246 12/07/06 08:13 Makis Petrat Shift Summary
  • 190243 12/07/06 07:24 Elena Kh. spot++
  • 190241 12/07/06 07:19 Elena Kh. Dummy 20cm
  • 190237 12/07/06 05:28 Elena Kh. TOF for runs #1365, 1366, 1367
  • 190236 12/07/06 05:23 Elena Kh. Target screens
  • 190235 12/07/06 05:21 Elena Kh. Target screens
  • 190230 12/07/06 04:02 Makis Petrat DVCS device HV alarm
  • 190223 12/07/06 00:01 Seonho Choi Swing shift summary (Dec. 6, 2006)
  • Wednesday

  • 190217 12/06/06 21:35 Kumbartzki Target charts
  • 190216 12/06/06 21:31 Kumbartzki Target screens
  • 190210 12/06/06 18:01 atrig Trigger-Download
  • 190205 12/06/06 15:59 Ole Hansen Shift Summary
  • 190202 12/06/06 15:04 Ole Hansen Hall A Tools & beam current
  • 190201 12/06/06 14:51 A. Camsonne Target increased limit output temperature
  • 190200 12/06/06 14:46 A. Camsonne Target He3 at 50 uA
  • 190196 12/06/06 14:37 A. Camsonne Beam back at 40 uA
  • 190193 12/06/06 12:58 A. Camsonne Target screens
  • 190192 12/06/06 12:57 A. Camsonne Target power heater current readback
  • 190189 12/06/06 12:08 Jack Gas System maintenance
  • 190188 12/06/06 11:30 Jack Carbon Dioxide supply
  • 190187 12/06/06 11:02 A . Camsonne Target temperature set to 7.5 K
  • 190174 12/06/06 07:15 atrig Trigger-Download
  • 190170 12/06/06 05:12 Elena Kh. D.ctimel for run# 1347
  • 190169 12/06/06 05:07 Elena Kh. Wrong run# for last D.ctimel
  • 190168 12/06/06 05:01 Adaq D.ctimel for run# 1946
  • 190165 12/06/06 04:45 Elena Kh. He3 target
  • 190159 12/06/06 03:39 Elena Kh. Dummy target
  • 190157 12/06/06 03:12 Adaq D.ctimel for run# 1345
  • 190151 12/06/06 00:17 atrig Trigger-Download
  • 190148 12/06/06 00:03 Seonho Choi Swing shift summary (Dec. 5, 2006)
  • Tuesday

  • 190146 12/05/06 23:58 atrig Trigger-Download
  • 190142 12/05/06 21:41 R. Michaels turned on end-of-run msg
  • 190140 12/05/06 21:35 G.M. Urciuol Target moved to He3
  • 190138 12/05/06 21:17 Jack Carbon dioxide supply
  • 190136 12/05/06 20:24 G.M. Urciuol Target moved to Dummy
  • 190132 12/05/06 18:10 Seonho Choi Kinematics change to #2
  • 190130 12/05/06 17:56 atrig Trigger-Download
  • 190129 12/05/06 17:40 G.M. Urciuol Target moved to He3 position
  • 190128 12/05/06 17:27 atrig Trigger-Download
  • 190126 12/05/06 16:54 A. Camsonne D.bit3 timing spectrum
  • 190125 12/05/06 16:47 G.M. Urciuol Target moved to dummy
  • 190122 12/05/06 16:27 A. Camsonne Shift summary day
  • 190120 12/05/06 16:25 A. Camsonne Coincidence time Run 1336
  • 190119 12/05/06 16:24 A. Camsonne Replay for T5 with split files
  • 190118 12/05/06 15:58 A. Camsonne small current excursion
  • 190115 12/05/06 14:07 A. Camsonne Increased ps1 and ps3 to 10
  • 190110 12/05/06 13:07 A. Camsonne Gas cerenkov after rough calibration
  • 190107 12/05/06 12:58 Margaziotis Target screens
  • 190106 12/05/06 12:56 Margaziotis Target move to 3He
  • 190105 12/05/06 12:54 A. Camsonne Database coefficient for gas cerenkov updated
  • 190094 12/05/06 08:41 A. Camsonne Coincidence time
  • 190093 12/05/06 08:04 suleiman Owl Shift Summary
  • 190090 12/05/06 07:57 Elena Kh. Dummy 20cm
  • 190088 12/05/06 07:22 Elena Kh. Target screens
  • 190085 12/05/06 06:10 Adaq Rates at S1-R HV at 70%
  • 190083 12/05/06 05:51 Elena Kh. RS1 HV to 70% of its original value
  • 190081 12/05/06 05:09 Adaq S1-R ADCs and Coincidence Time at 0.60 HV
  • 190080 12/05/06 05:07 Adaq S1-R ADCs and Coincidence Time at 0.67 HV
  • 190079 12/05/06 05:04 Adaq S1-R ADCs and Coincidence Time at 0.75 HV
  • 190078 12/05/06 04:47 Adaq Rates at 0.75 HV S1-R
  • 190075 12/05/06 04:41 Adaq Back to 67% HV on S1-R
  • 190074 12/05/06 04:17 Adaq Rates at 0.75 HV on S1-R
  • 190072 12/05/06 04:04 Adaq Reduce S1 Right HV to 0.75 of its value
  • 190064 12/05/06 01:44 Adaq Rates at 67% of original S1 right HV
  • 190062 12/05/06 01:42 Adaq S1 Right HV Reduced to 67% of their original value
  • 190060 12/05/06 01:10 Adaq Scaler Rates after S1 Right HV was reduced to 60% of its original value
  • 190058 12/05/06 01:05 Elena Kh. New RS1 HV
  • 190057 12/05/06 00:59 Elena Kh. RS1 HV before to change them
  • 190055 12/05/06 00:38 Adaq Right Arm HV at the Start of Owl Shift
  • 190054 12/05/06 00:34 Adaq Target Screens
  • 190053 12/05/06 00:32 Adaq Hall A General Tool
  • 190052 12/05/06 00:29 Adaq spot++
  • Monday

  • 190048 12/04/06 23:57 Seonho Choi Swing Shift Summary (Mon. Dec. 4, 2006)
  • 190047 12/04/06 23:08 A. Camsonne Replay with cut on T5
  • 190044 12/04/06 22:47 A. Camsonne Coincidence time on He3
  • 190042 12/04/06 21:23 folts tech VDC covers
  • 190040 12/04/06 20:54 Margaziotis More target screens
  • 190039 12/04/06 20:47 Margaziotis Target screens
  • 190037 12/04/06 20:36 atrig Trigger-Download
  • 190034 12/04/06 20:34 A. Camsonne Gas Cerenkov PMT 2 reduced to 1600 V
  • 190032 12/04/06 18:53 Seonho Choi Strange T1 and T3 rates - call for help
  • 190026 12/04/06 17:40 Seonho Choi BLM problem
  • 190025 12/04/06 16:50 jpchen/dimit set target operating temperatue to 7.0K
  • 190023 12/04/06 16:21 jpchen Day shift summary
  • 190022 12/04/06 16:13 jpchen ESR problem, CHL and target OK
  • 190021 12/04/06 16:09 jpchen ESR problem, Left Q1/dipole and right Q1/Q3 are gone.
  • 190020 12/04/06 15:45 Seonho Choi Swing Shift Summary (Dec. 3, 2006)
  • 190018 12/04/06 15:34 JPChen spot++
  • 190016 12/04/06 15:24 atrig Trigger-Download
  • 190013 12/04/06 14:55 G.M. Urciuol Target moved
  • 190012 12/04/06 14:38 jpchen S1_R HV restored back to normal
  • 190007 12/04/06 13:59 jpchen lowered S1_R HV by another 8% for test runs on dummy
  • 190006 12/04/06 13:57 jpchen lowered S1_R HV by another 8%
  • 190004 12/04/06 13:44 jpchen He3 target copration condition
  • 190003 12/04/06 12:56 jpchen cryotarget switched from LH2 to He3
  • 190002 12/04/06 12:04 bogdanw Left arm Gas Cerenkov
  • 190001 12/04/06 11:19 jpchen Left gas Cherenkov PMT HV changes
  • 189999 12/04/06 09:32 jpchen L_Cherenkov PMT#2 TDC spectrun
  • 189998 12/04/06 09:17 jpchen HV setting with S1_R lowered by 8%
  • 189997 12/04/06 09:11 jpchen Hall A in controlled access for target switch (LH2->He3)
  • 189993 12/04/06 08:16 jpchen/Elena S1_R HV lowered by 8%
  • 189991 12/04/06 08:08 Makis Petrat Shift Summary
  • 189985 12/04/06 07:03 atrig Trigger-Download
  • 189978 12/04/06 04:11 atrig Trigger-Download
  • 189977 12/04/06 04:10 atrig Trigger-Download
  • Sunday

  • 189962 12/03/06 23:13 Margaziotis Target screens
  • 189960 12/03/06 22:29 atrig Trigger-Download
  • 189941 12/03/06 19:47 atrig Trigger-Download
  • 189929 12/03/06 17:19 Seonho Choi Swing Shift Summary (Dec. 3, 2006)
  • 189919 12/03/06 16:04 G.M. Urciuol 12/3/2006 Day shift report
  • 189915 12/03/06 15:51 atrig Trigger-Download
  • 189914 12/03/06 15:46 atrig Trigger-Download
  • 189913 12/03/06 15:45 atrig Trigger-Download
  • 189909 12/03/06 14:25 P. Decowski wrong target name in runs 1265 and 1266
  • 189903 12/03/06 14:02 atrig Trigger-Download
  • 189896 12/03/06 12:44 atrig Trigger-Download
  • 189892 12/03/06 11:50 R. Michaels Coinc. trigger checks
  • 189891 12/03/06 11:44 P. Decowski Target screens 2
  • 189890 12/03/06 11:41 P. Decowski Target screens
  • 189880 12/03/06 11:13 folts tech/bogies
  • 189879 12/03/06 11:10 atrig Trigger-Download
  • 189878 12/03/06 08:57 atrig Trigger-Download
  • 189876 12/03/06 08:19 suleiman Owl Shift Summary
  • 189865 12/03/06 04:12 suleiman Reboot the correct IOC this time but still the problem was not fixed
  • 189862 12/03/06 03:54 Adaq Target Screens
  • 189861 12/03/06 03:49 Adaq Shift Screens
  • 189860 12/03/06 03:48 Adaq Spot
  • 189857 12/03/06 02:36 Adaq Coda Crashed - ROC1 reboot
  • 189855 12/03/06 02:32 Adaq Moved to 20 cm Dummy
  • 189851 12/03/06 01:54 suleiman Right Arn Bogies are not changing status
  • 189844 12/03/06 01:37 Elena Kh. Kinematics #8
  • 189838 12/03/06 00:37 Elena Kh. Kinematics 7: angles.
  • 189836 12/03/06 00:26 atrig Trigger-Download
  • 189835 12/03/06 00:10 G.M. Urciuol 12/2/20006 swing shift report
  • Saturday

  • 189828 12/02/06 22:00 atrig Trigger-Download
  • 189822 12/02/06 20:55 folts target camera
  • 189820 12/02/06 20:37 Seonho Choi Run 1222 and 1223 replay finished
  • 189819 12/02/06 20:37 Seonho Choi Run 1222 and 1223 replay finished
  • 189814 12/02/06 19:55 atrig Trigger-Download
  • 189807 12/02/06 18:50 atrig Trigger-Download
  • 189802 12/02/06 17:37 Seonho Choi Right bogies fixed (thanks to Ed)
  • 189801 12/02/06 17:27 atrig Trigger-Download
  • 189800 12/02/06 16:56 Seonho Choi Calling for a controlled access to fix right bogies problem
  • 189799 12/02/06 16:50 Seonho Choi Right bogies communication problem
  • 189797 12/02/06 16:05 bogdanw shift summary
  • 189795 12/02/06 15:38 Margaziotis Target screens
  • 189794 12/02/06 15:35 Margaziotis Target moved to LH2
  • 189790 12/02/06 15:03 atrig Trigger-Download
  • 189788 12/02/06 14:43 R. Michaels need to check TC calib
  • 189783 12/02/06 12:59 R. Michaels Added a coinc. time diff histo
  • 189782 12/02/06 12:45 Margaziotis Target to Carbon
  • 189780 12/02/06 11:43 atrig Trigger-Download
  • 189779 12/02/06 11:41 bogdanw kinematic "ep 2"
  • 189776 12/02/06 10:30 Margaziotis Target to Dummy
  • 189772 12/02/06 09:44 Adaq S2 Left TDCs now working
  • 189768 12/02/06 07:53 suleiman Owl Shift Summary
  • 189767 12/02/06 07:28 Adaq Spot
  • 189766 12/02/06 07:26 Adaq Hall A General Tools
  • 189764 12/02/06 06:45 atrig Trigger-Download
  • 189763 12/02/06 06:36 A. Camsonne Left arm S2 few counts
  • 189759 12/02/06 05:26 A. Camsonne HV change S2m
  • 189756 12/02/06 04:44 suleiman New S1 Right ADCs
  • 189753 12/02/06 04:05 Elena Kh. Target screens
  • 189741 12/02/06 02:57 Adaq Changing few HV on S1 Right
  • 189739 12/02/06 02:51 Adaq Turn OFF Cerenkov HV PMT 5
  • 189732 12/02/06 01:45 atrig Trigger-Download
  • 189729 12/02/06 01:21 atrig Trigger-Download
  • 189726 12/02/06 00:57 atrig Trigger-Download
  • 189724 12/02/06 00:39 Adaq Hall A General Tools Screen
  • 189723 12/02/06 00:37 A. Camsonne Spectrometer kin change
  • 189721 12/02/06 00:10 Adaq Insertable HWP is IN at 12/01/06 16:54
  • 189719 12/02/06 00:00 Seonho Choi Swing shift summary (Dec. 1, 2006)
  • Friday

  • 189717 12/01/06 23:55 atrig Trigger-Download
  • 189714 12/01/06 23:28 atrig Trigger-Download
  • 189711 12/01/06 22:52 atrig Trigger-Download
  • 189710 12/01/06 22:42 atrig Trigger-Download
  • 189707 12/01/06 22:06 atrig Trigger-Download
  • 189705 12/01/06 21:48 P. Decowski Target screen 1 once more
  • 189703 12/01/06 21:41 atrig Trigger-Download
  • 189702 12/01/06 21:39 P. Decowski Target screens 2
  • 189700 12/01/06 21:36 P. Decowski Target screens 1
  • 189698 12/01/06 21:13 Seonho Choi RF problem fixed. Beam restoration.
  • 189697 12/01/06 20:27 Seonho Choi RF problem with machine
  • 189696 12/01/06 20:10 atrig Trigger-Download
  • 189691 12/01/06 19:43 atrig Trigger-Download
  • 189685 12/01/06 17:47 Seonho Choi BPM Strip Chart
  • 189684 12/01/06 17:46 Seonho Choi General Tools snapshot
  • 189683 12/01/06 17:42 Seonho Choi HRSe NMR lock
  • 189679 12/01/06 17:12 Seonho Choi HRSh (right) spectro back to elastic
  • 189678 12/01/06 17:06 Seonho Choi HRSe (left) momentum change
  • 189676 12/01/06 16:26 Seonho Choi Beam back
  • 189673 12/01/06 16:01 LeRose Right Arm at +2%
  • 189672 12/01/06 16:00 Saha Shift Summary
  • 189671 12/01/06 15:50 LeRose Rebooting Right NMR
  • 189670 12/01/06 15:43 LeRose More Target Screens
  • 189669 12/01/06 15:29 LeRose Target screens
  • 189662 12/01/06 10:42 Elena Kh. run# 1178: R_P=1.951
  • 189661 12/01/06 10:32 LeRose Right Arm -2% Change/Q3 positive
  • 189657 12/01/06 09:32 LeRose Right Q3 Negative
  • 189650 12/01/06 08:02 suleiman Owl Shift Summary
  • 189649 12/01/06 07:37 Adaq Right Arm back to nominal
  • 189638 12/01/06 05:19 Adaq Changed Left to 0%, Cycled Q2/3
  • 189634 12/01/06 04:46 Adaq Changed Left by -3%
  • 189631 12/01/06 04:13 Adaq Changed Left by +3%
  • 189629 12/01/06 04:06 Adaq Target Screens
  • 189628 12/01/06 03:52 Adaq Moved to H2
  • 189623 12/01/06 03:22 Adaq Moved target to Optics
  • 189619 12/01/06 02:30 Adaq Moved target to Empty
  • 189616 12/01/06 02:09 Adaq Move Target to 20 cm Dummy
  • 189615 12/01/06 02:01 suleiman Hall C Energy Modulation
  • 189611 12/01/06 00:59 Adaq Spot Plot
  • 189610 12/01/06 00:57 Adaq Target Screens
  • 189607 12/01/06 00:50 Adaq Hall A General Tools
  • 189606 12/01/06 00:48 atrig Trigger-Download